graph the function of y= 5-4 cos x/2
HELP! TY@

Answer by lwsshak3(11628)

graph the function of y= 5-4 cos x/2
rewrite equation:
y=-4 cos x/2+5
Standard form of equation for cos function: y=Acos(Bx-C), A=amplitude, Period=2π/B, Phase shift=C/B
For given function:y=-4 cos x/2+5
Amplitude=4
B=1/2
Period=2π/B=2π/(1/2)=4π
1/4 period=π
C=0
Phase shift: none
curve is bumped up 5 units
..
Graphing:
I don't have the means to graph the function, but I will develop the coordinates for one period with which you can use to draw the graph yourself.
Start with the basic cos function:
(0,1), (π,0), (2π,-1), (3π,0), (0,1)
reflect about the x-axis:
(0,-1), (π,0), (2π,1), (3π,0), (0,-1)
raise amplitude to 4 units
(0,-4), (π,0), (2π,4), (3π,0), (0,-4)
finally, bump up curve 5 units
(0,1), (π,5), (2π,9), (3π,5), (0,1) (final configuration)